# CondTSF: One-line Plugin of Dataset Condensation for Time Series Forecasting

### v3

@@ Abstract
Dataset condensation is a newborn technique that generates a small dataset that can be used in training deep neural networks (DNNs) to lower storage and training costs. The objective of dataset condensation is to ensure that the model trained with the synthetic dataset can perform comparably to the model trained with full datasets. However, existing methods predominantly concentrate on classification tasks, posing challenges in their adaptation to time series forecasting (TS-forecasting). This challenge arises from disparities in the evaluation of synthetic data. In classification, the synthetic data is considered well-distilled if the model trained with the full dataset and the model trained with the synthetic dataset yield identical labels for the same input, regardless of variations in output logits distribution. Conversely, in TS-forecasting, the effectiveness of synthetic data distillation is determined by the distance between predictions of the two models. The synthetic data is deemed well-distilled only when all data points within the predictions are similar. Consequently, TS-forecasting has a more rigorous evaluation methodology compared to classification. To mitigate this gap, we theoretically analyze the optimization objective of dataset condensation for TS-forecasting and propose a new one-line plugin of dataset condensation for TS-forecasting designated as Dataset Condensation for Time Series Forecasting (CondTSF) based on our analysis. Plugging CondTSF into previous dataset condensation methods facilitates a reduction in the distance between the predictions of the model trained with the full dataset and the model trained with the synthetic dataset, thereby enhancing performance. We conduct extensive experiments on eight commonly used time series datasets. CondTSF consistently improves the performance of all previous dataset condensation methods across all datasets, particularly at low condensing ratios.1
@@ 3 Preliminaries
Dataset Condensation for TS-forecasting Given a time series dataset, we split the dataset into a train set and a test set. In this paper, we denote the train set as $\boldsymbol{f}$ and the test set as $\boldsymbol{x}$. We denote the synthetic dataset as $\boldsymbol{s}$. The synthetic dataset $\boldsymbol{s}$ is a small dataset distilled from the full train set $\boldsymbol{f}$. We employ $\mathcal{M}_{\theta}$ as a neural network parameterized by $\theta$. Without losing generality, we suppose the model $\mathcal{M}_{\theta}$ is using historical sequence $\boldsymbol{x}_{t:t+m}$ with length $m$ to predict future sequence $\boldsymbol{x}_{t+m:t+m+n}$ with length $n$. Given the test set $\boldsymbol{x}$, we formulate the test error of $\mathcal{M}_{\theta}$ as the error between the prediction of $\mathcal{M}_{\theta}$ on test input $\boldsymbol{x}_{t:t+m}$ and the test label $\boldsymbol{x}_{t+m:t+m+n}$, as shown in Eq.1.
@@ 4.1 Decomposition
Then we decompose the upper bound of $\mathcal{L}_{test}(\mathcal{M}_{\theta_{s,test}},\boldsymbol{x})$ into two terms, as shown in Thm.1.
@@
To prove Thm.1, we use linear models for further analysis since linear models can be both effective and efficient in TS-forecasting[45]. Given a linear model $\mathcal{M}_{\theta}(\boldsymbol{x})=\theta\boldsymbol{x}$, its second and higher order gradient is zero. Therefore first-order Taylor Expansion is sufficient to obtain the accurate prediction of the model. We prove Thm.1 by applying the property of the first-order Taylor Expansion and triangular inequality of norm functions. The complete proof is in App.A.1. Hence we decompose the optimization objective of dataset condensation for TS-forecasting into two optimizable terms, namely value term and gradient term. For value term, it ensures $\mathcal{M}_{\theta_{s,test}}$ and $\mathcal{M}_{\theta_{f,test}}$ are similar in prediction values. For gradient term, it ensures the predictions of $\mathcal{M}_{\theta_{s,test}}$ and $\mathcal{M}_{\theta_{f,test}}$ are similar in gradient. Optimizing these two terms can optimize the upper bound of the original optimization objective, and therefore indirectly optimize the original optimization objective in Eq.3.
@@ 4.4 CondTSF
In this section, we develop a one-line plugin called CondTSF to minimize the label error $\mathcal{L}_{label}$ in Eq.11 so that the value term can be optimized. CondTSF is a lightweight one-line plugin, no backpropagation or gradient is required during the update. CondTSF utilizes a simple yet effective additive method to iteratively update the synthetic data $\boldsymbol{s}$ and minimize the label error $\mathcal{L}_{label}$. In the $i_{\text{th}}$ update iteration, CondTSF uses the prediction of expert model $\mathcal{M}_{f,test}(\boldsymbol{s}_{t^{\prime}:t^{\prime}+m})$ to update synthetic label $\boldsymbol{s}_{t^{\prime}+m:t^{\prime}+m+n}$. The update process is shown in Eq.12.
@@
Since the update ratio has a value of $10<\beta<$ 1, CondTSF lowers the label error $\mathcal{L}_{label}$ exponentially in each update iteration and solves the optimization problem for the value term. As a plugin module, CondTSF is used to update once for every $G$ iterations of parameter matching methods. In this way, the gradient term and the value term can be optimized synchronously. The algorithm is shown in Alg.1. We also formulate the complete condensation process using CondTSF, as shown in Fig.2.
@@ 5.1 Experiment Settings
Model Settings: We plug CondTSF into existing dataset condensation models based on parameter matching, including DC[51], MTT[3], PP[21], TESLA[5], FTD[7] and DATM[11] to prove the effectiveness of CondTSF. We also conduct experiments on non-parameter-matching based methods, including DM[50], IDM[52], KIP[33], FRePo[55] to prove that optimizing value term only also helps boost the performance. The experiment setting and results are shown in App.C. We use DLinear[45] as the expert model to perform dataset condensation since DLinear is a linear model.
@@ 6 Limitations
The limitation of this work is that we use linear models in our analysis so that the gradient of a model on input is the parameter of the model. Therefore, only linear models like DLinear[45] are solid enough to be an expert model for dataset condensation. The analysis no longer holds when it comes to more complicated models. However, experiments show that CondTSF is also effective with non-parameter-matching methods and non-linear models, which merits further exploration.
@@ A.1 Complete Proof for Theorem 1
In Eq.15, we prove that minimizing the distance between $\mathcal{M}_{\theta_{s,test}}(\boldsymbol{x}_{t:t+m})$ and $\mathcal{M}_{\theta_{f,test}}(\boldsymbol{x}_{t:t+m})$ is equivalent to minimizing the upper bound of the original optimization objective. Then we decompose the distance between predictions of $\mathcal{M}_{\theta_{s,test}}$ and $\mathcal{M}_{\theta_{f,test}}$ into two optimizable terms for better optimization. We use linear models for further analysis since linear models can be both effective and efficient in TS-forecasting tasks[45]. Given a linear model $\mathcal{M}_{\theta}(\boldsymbol{x})=\theta\boldsymbol{x}$, its second and higher order gradient is zero, i.e. $\nabla^{k}\mathcal{M}_{\theta}(\boldsymbol{x})=\boldsymbol{0},\forall k\geq 2$. Therefore, first-order Taylor Expansion can be utilized to get the prediction of the model $\mathcal{M}_{\theta}$ on test data $\boldsymbol{x}_{t:t+m}$ using the prediction and gradient of the model $\mathcal{M}_{\theta}$ on arbitrary synthetic data $\boldsymbol{s}_{t^{\prime}:t^{\prime}+m}$. The process is formulated in Eq.16.
@@ B.1 Standard Ratio Condensation
Appendix C Performance of Non-parameter-matching Based Methods
@@
Appendix D Ablation Study of CondTSF

### v4

@@ Abstract
Dataset condensation is a newborn technique that generates a small dataset that can be used in training deep neural networks (DNNs) to lower storage and training costs. The objective of dataset condensation is to ensure that the model trained with the synthetic dataset can perform comparably to the model trained with full datasets. However, existing methods predominantly concentrate on classification tasks, posing challenges in their adaptation to time series forecasting (TS-forecasting). This challenge arises from disparities in the evaluation of synthetic data. In classification, the synthetic data is considered well-distilled if the model trained with the full dataset and the model trained with the synthetic dataset yield identical labels for the same input, regardless of variations in output logits distribution. Conversely, in TS-forecasting, the effectiveness of synthetic data distillation is determined by the distance between predictions of the two models. The synthetic data is deemed well-distilled only when all data points within the predictions are similar. Consequently, TS-forecasting has a more rigorous evaluation methodology compared to classification. To mitigate this gap, we theoretically analyze the optimization objective of dataset condensation for TS-forecasting and propose a new one-line plugin of dataset condensation for TS-forecasting designated as Dataset Condensation for Time Series Forecasting (CondTSF) based on our analysis. Plugging CondTSF into previous dataset condensation methods facilitates a reduction in the distance between the predictions of the model trained with the full dataset and the model trained with the synthetic dataset, thereby enhancing performance. We conduct extensive experiments on eight commonly used time series datasets. CondTSF consistently improves the performance of all previous dataset condensation methods across all datasets, particularly at low condensing ratios.
@@ 3 Preliminaries
Dataset Condensation for TS-forecasting Given a time series dataset, we split the dataset into a train set and a test set. In this paper, we denote the train set as $\boldsymbol{f}$ and the test set as $\boldsymbol{x}$. We denote the synthetic dataset as $\boldsymbol{s}$. The synthetic dataset $\boldsymbol{s}$ is a small dataset distilled from the full train set $\boldsymbol{f}$. Train set $\boldsymbol{f}$, test set $\boldsymbol{x}$, and synthetic dataset $\boldsymbol{s}$ are all vectors. We employ $\mathcal{M}_{\theta}$ as a neural network parameterized by $\theta$. Without losing generality, we suppose the model $\mathcal{M}_{\theta}$ is using historical sequence $\boldsymbol{x}_{t:t+m}$ with length $m$ to predict future sequence $\boldsymbol{x}_{t+m:t+m+n}$ with length $n$. Given the test set $\boldsymbol{x}$, we formulate the test error of $\mathcal{M}_{\theta}$ as the error between the prediction of $\mathcal{M}_{\theta}$ on test input $\boldsymbol{x}_{t:t+m}$ and the test label $\boldsymbol{x}_{t+m:t+m+n}$, as shown in Eq.1.
@@ 4.1 Decomposition
Then we decompose the upper bound of $\mathcal{L}_{test}(\mathcal{M}_{\theta_{s,test}},\boldsymbol{x})$ into two terms, as shown in Thm.1. We utilize Taylor Expansion in the proof of Thm.1. For each real test data $x_{t:t+m}$, we can arbitrarily choose position $t^{\prime}$ and get synthetic data $s_{t^{\prime}:t^{\prime}+m}$. Then we can perform Taylor Expansion with $s_{t^{\prime}:t^{\prime}+m}$ to obtain the value of $\mathcal{M}_{\theta_{s,test}}(\boldsymbol{x}_{t:t+m})$ and $\mathcal{M}_{\theta_{f,test}}(\boldsymbol{x}_{t:t+m})$.
@@
To prove Thm.1, we use linear models for further analysis since linear models can be both effective and efficient in TS-forecasting[45]. Given a linear model $\mathcal{M}_{\theta}(\boldsymbol{x})=\theta\boldsymbol{x}$, its second and higher order gradient is zero. Therefore first-order Taylor Expansion is sufficient to obtain the accurate prediction of the model. Meanwhile, if $\mathcal{M}_{\theta}$ is a non-linear model, we ignore the higher-order terms of Taylor Expansion. We prove Thm.1 by applying the property of the first-order Taylor Expansion and triangular inequality of norm functions. The complete proof is in App.A.1. Hence we decompose the optimization objective of dataset condensation for TS-forecasting into two optimizable terms, namely value term and gradient term. For value term, it ensures $\mathcal{M}_{\theta_{s,test}}$ and $\mathcal{M}_{\theta_{f,test}}$ are similar in prediction values. For gradient term, it ensures the predictions of $\mathcal{M}_{\theta_{s,test}}$ and $\mathcal{M}_{\theta_{f,test}}$ are similar in gradient. Optimizing these two terms can optimize the upper bound of the original optimization objective, and therefore indirectly optimize the original optimization objective in Eq.3.
@@ 4.4 CondTSF
In this section, we develop a one-line plugin called CondTSF to minimize the label error $\mathcal{L}_{label}$ in Eq.11 so that the value term can be optimized. CondTSF is a lightweight one-line plugin, no backpropagation or gradient is required during the update. CondTSF utilizes a simple yet effective additive method to iteratively update the synthetic data $\boldsymbol{s}$ and minimize the label error $\mathcal{L}_{label}$. In TS-forecasting, when generating training data, the data is usually sampled overlap from the dataset. Inspired by the overlap property, we utilize an additive method in CondTSF to gradually update the synthetic data to avoid vibrations. In the $i_{\text{th}}$ update iteration, CondTSF uses the prediction of expert model $\mathcal{M}_{f,test}(\boldsymbol{s}_{t^{\prime}:t^{\prime}+m})$ to update synthetic label $\boldsymbol{s}_{t^{\prime}+m:t^{\prime}+m+n}$. The update process is shown in Eq.12.
@@
Since the update ratio has a value of $10<\beta<$ 1, CondTSF lowers the label error $\mathcal{L}_{label}$ exponentially in each update iteration and solves the optimization problem for the value term. As a plugin module, CondTSF is used to update once for every $G$ iteration of parameter matching methods. In this way, the gradient term and the value term can be optimized synchronously. The algorithm is shown in Alg.1. We also formulate the complete condensation process using CondTSF, as shown in Fig.2.
@@ 5.1 Experiment Settings
Model Settings: We plug CondTSF into existing dataset condensation models based on parameter matching, including DC[51], MTT[3], PP[21], TESLA[5], FTD[7] and DATM[11] to prove the effectiveness of CondTSF. We also conduct experiments on non-parameter-matching based methods, including DM[50], IDM[52], KIP[33], FRePo[55] to prove that optimizing value term only also helps boost the performance. The experiment setting and results are shown in App.E. We use DLinear[45] as the expert model to perform dataset condensation since DLinear is a linear model.
@@ 6 Limitations
The limitation of this work is that we use linear models in our analysis so that the gradient of a model on input is the parameter of the model. Therefore, only linear models like DLinear[45] are solid enough to be an expert model for dataset condensation. The analysis no longer holds when it comes to more complicated models. However, experiments in App.D and App.E show that CondTSF is also effective with non-parameter-matching methods and non-linear models, which merits further exploration.
@@ 7 Conclusion
Acknowledgements
This work was sponsored by National Natural Science Foundation of China under Grant No. 62102246, 62272301, and Provincial Key Research and Development Program of Zhejiang under Grant No. 2021C01034. Part of the work was done when the students were doing internships at Yunqi Academy of Engineering.
@@ A.1 Complete Proof for Theorem 1
In Eq.15, we prove that minimizing the distance between $\mathcal{M}_{\theta_{s,test}}(\boldsymbol{x}_{t:t+m})$ and $\mathcal{M}_{\theta_{f,test}}(\boldsymbol{x}_{t:t+m})$ is equivalent to minimizing the upper bound of the original optimization objective. Then we decompose the distance between predictions of $\mathcal{M}_{\theta_{s,test}}$ and $\mathcal{M}_{\theta_{f,test}}$ into two optimizable terms for better optimization. We use linear models for further analysis since linear models can be both effective and efficient in TS-forecasting tasks[45]. Given a linear model $\mathcal{M}_{\theta}(\boldsymbol{x})=\theta\boldsymbol{x}$, its second and higher order gradient is zero, i.e. $\nabla^{k}\mathcal{M}_{\theta}(\boldsymbol{x})=\boldsymbol{0},\forall k\geq 2$. Therefore, first-order Taylor Expansion can be utilized to get the prediction of the model $\mathcal{M}_{\theta}$ on test data $\boldsymbol{x}_{t:t+m}$ using the prediction and gradient of the model $\mathcal{M}_{\theta}$ on arbitrary synthetic data $\boldsymbol{s}_{t^{\prime}:t^{\prime}+m}$. The process is formulated in Eq.16. Meanwhile, it is worth mentioning that if $\mathcal{M}_{\theta}$ is a non-linear model, then the second and higher-order terms of the Taylor Expansion are ignored in Eq.16.
@@ B.1 Standard Ratio Condensation
Appendix C Performance of CondTSF with Non-parameter-matching Based Methods
@@
Appendix D Performance of CondTSF with Non-Linear Expert Models
We distill the dataset using the standard condensing ratio. The information on condensation is shown in Table.4. We conduct experiments on distilling dataset with non-linear expert models. We use MTT[3], TESLA[5], and DATM[11] as backbone methods. The performance of using a CNN as the expert model is shown in Table.9, and the performance of using a 3-layer-MLP as the expert model is shown in Table.10.
Results show that CondTSF is also effective with non-linear expert models.
Table 9: Distill performance of different dataset condensation methods with CNN as the expert model. For each method, ✗means CondTSF is not used, ✓means CondTSF is used, and $\downarrow$ means the decreased percentage of test error after CondTSF is applied. Five synthetic datasets are distilled and the average and standard deviation are reported.
CondTSF ExchangeRate Weather Electricity Traffic
MAE MSE MAE MSE MAE MSE MAE MSE
Random - 0.830$\pm$0.059 1.002$\pm$0.127 0.504$\pm$0.018 0.526$\pm$0.030 0.762$\pm$0.016 0.882$\pm$0.034 0.732$\pm$0.023 1.041$\pm$0.053
MTT 0.372$\pm$0.028 0.237$\pm$0.031 0.314$\pm$0.013 0.278$\pm$0.007 0.482$\pm$0.009 0.393$\pm$0.012 0.662$\pm$0.011 0.906$\pm$0.026
0.140$\pm$0.011 0.063$\pm$0.003 0.246$\pm$0.008 0.120$\pm$0.004 0.357$\pm$0.005 0.267$\pm$0.005 0.451$\pm$0.013 0.519$\pm$0.009
$\downarrow$ 62.4% 73.4% 21.7% 56.8% 25.9% 32.1% 31.9% 42.7%
TESLA 0.378$\pm$0.007 0.242$\pm$0.007 0.310$\pm$0.015 0.292$\pm$0.012 0.516$\pm$0.005 0.430$\pm$0.009 0.655$\pm$0.020 0.900$\pm$0.045
0.134$\pm$0.012 0.058$\pm$0.002 0.253$\pm$0.007 0.137$\pm$0.002 0.374$\pm$0.008 0.267$\pm$0.007 0.528$\pm$0.011 0.632$\pm$0.020
$\downarrow$ 64.6% 76.0% 18.4% 53.1% 27.5% 37.9% 19.4% 29.8%
DATM 0.331$\pm$0.011 0.179$\pm$0.013 0.335$\pm$0.009 0.294$\pm$0.008 0.504$\pm$0.008 0.432$\pm$0.009 0.587$\pm$0.009 0.742$\pm$0.014
0.137$\pm$0.017 0.059$\pm$0.004 0.291$\pm$0.005 0.261$\pm$0.004 0.355$\pm$0.006 0.252$\pm$0.006 0.452$\pm$0.009 0.526$\pm$0.011
$\downarrow$ 58.6% 67.0% 13.1% 11.2% 29.6% 41.7% 23.0% 29.1%
CondTSF ETTm1 ETTm2 ETTh1 ETTh2
MAE MSE MAE MSE MAE MSE MAE MSE
Random - 0.691$\pm$0.012 0.856$\pm$0.030 0.723$\pm$0.027 0.846$\pm$0.052 0.727$\pm$0.015 0.931$\pm$0.030 0.722$\pm$0.015 0.847$\pm$0.033
MTT 0.550$\pm$0.006 0.585$\pm$0.011 0.347$\pm$0.008 0.205$\pm$0.009 0.644$\pm$0.013 0.788$\pm$0.036 0.371$\pm$0.016 0.245$\pm$0.020
0.482$\pm$0.007 0.507$\pm$0.007 0.236$\pm$0.009 0.111$\pm$0.005 0.460$\pm$0.008 0.437$\pm$0.014 0.297$\pm$0.011 0.173$\pm$0.009
$\downarrow$ 12.4% 13.3% 32.0% 45.9% 28.6% 44.5% 19.9% 29.4%
TESLA 0.544$\pm$0.003 0.583$\pm$0.007 0.359$\pm$0.013 0.222$\pm$0.016 0.634$\pm$0.010 0.757$\pm$0.037 0.365$\pm$0.008 0.240$\pm$0.009
0.499$\pm$0.003 0.492$\pm$0.007 0.251$\pm$0.009 0.119$\pm$0.005 0.473$\pm$0.009 0.458$\pm$0.018 0.293$\pm$0.004 0.170$\pm$0.002
$\downarrow$ 8.3% 15.6% 30.1% 46.4% 25.4% 39.5% 19.7% 29.2%
DATM 0.566$\pm$0.008 0.598$\pm$0.011 0.318$\pm$0.013 0.174$\pm$0.014 0.633$\pm$0.009 0.773$\pm$0.025 0.349$\pm$0.012 0.221$\pm$0.012
0.518$\pm$0.007 0.521$\pm$0.012 0.231$\pm$0.002 0.107$\pm$0.001 0.451$\pm$0.005 0.418$\pm$0.007 0.290$\pm$0.007 0.165$\pm$0.005
$\downarrow$ 8.5% 12.9% 27.4% 38.5% 28.8% 45.9% 16.9% 25.3%
Table 10: Distill performance of different dataset condensation methods with 3-layer-MLP as the expert model. For each method, ✗means CondTSF is not used, ✓means CondTSF is used, and $\downarrow$ means the decreased percentage of test error after CondTSF is applied. Five synthetic datasets are distilled and the average and standard deviation are reported.
CondTSF ExchangeRate Weather Electricity Traffic
MAE MSE MAE MSE MAE MSE MAE MSE
Random - 0.932$\pm$0.058 1.243$\pm$0.144 0.562$\pm$0.013 0.642$\pm$0.030 0.796$\pm$0.013 0.955$\pm$0.036 0.751$\pm$0.012 1.112$\pm$0.025
MTT 0.364$\pm$0.027 0.211$\pm$0.030 0.311$\pm$0.007 0.297$\pm$0.009 0.475$\pm$0.011 0.380$\pm$0.016 0.633$\pm$0.011 0.841$\pm$0.021
0.139$\pm$0.004 0.034$\pm$0.002 0.248$\pm$0.015 0.135$\pm$0.006 0.375$\pm$0.007 0.268$\pm$0.008 0.501$\pm$0.008 0.589$\pm$0.016
$\downarrow$ 61.8% 83.9% 20.3% 54.5% 21.1% 29.5% 20.9% 30.0%
TESLA 0.352$\pm$0.012 0.209$\pm$0.010 0.297$\pm$0.001 0.272$\pm$0.003 0.525$\pm$0.004 0.462$\pm$0.008 0.594$\pm$0.009 0.751$\pm$0.022
0.128$\pm$0.017 0.027$\pm$0.007 0.252$\pm$0.006 0.135$\pm$0.001 0.397$\pm$0.012 0.297$\pm$0.014 0.488$\pm$0.006 0.577$\pm$0.009
$\downarrow$ 63.6% 87.1% 15.2% 50.4% 24.4% 35.7% 17.8% 23.2%
DATM 0.326$\pm$0.016 0.177$\pm$0.016 0.349$\pm$0.002 0.301$\pm$0.001 0.517$\pm$0.008 0.441$\pm$0.011 0.622$\pm$0.007 0.785$\pm$0.005
0.141$\pm$0.005 0.042$\pm$0.001 0.254$\pm$0.010 0.126$\pm$0.004 0.385$\pm$0.009 0.280$\pm$0.008 0.496$\pm$0.006 0.582$\pm$0.011
$\downarrow$ 56.7% 76.3% 27.2% 58.1% 25.5% 36.5% 20.3% 25.9%
CondTSF ETTm1 ETTm2 ETTh1 ETTh2
MAE MSE MAE MSE MAE MSE MAE MSE
Random - 0.692$\pm$0.013 0.855$\pm$0.027 0.762$\pm$0.012 0.955$\pm$0.037 0.687$\pm$0.010 0.838$\pm$0.022 0.763$\pm$0.019 0.945$\pm$0.050
MTT 0.564$\pm$0.012 0.655$\pm$0.022 0.362$\pm$0.011 0.219$\pm$0.012 0.615$\pm$0.002 0.732$\pm$0.015 0.354$\pm$0.011 0.228$\pm$0.012
0.493$\pm$0.007 0.511$\pm$0.007 0.245$\pm$0.022 0.098$\pm$0.015 0.423$\pm$0.006 0.369$\pm$0.008 0.285$\pm$0.009 0.145$\pm$0.005
$\downarrow$ 12.6% 22.0% 32.3% 55.3% 31.2% 49.6% 19.5% 36.4%
TESLA 0.541$\pm$0.002 0.570$\pm$0.004 0.341$\pm$0.014 0.197$\pm$0.013 0.606$\pm$0.007 0.745$\pm$0.022 0.337$\pm$0.008 0.210$\pm$0.006
0.487$\pm$0.003 0.493$\pm$0.004 0.250$\pm$0.009 0.119$\pm$0.005 0.438$\pm$0.011 0.398$\pm$0.013 0.283$\pm$0.007 0.145$\pm$0.006
$\downarrow$ 10.0% 13.5% 26.7% 39.6% 27.7% 46.6% 16.0% 31.0%
DATM 0.558$\pm$0.009 0.611$\pm$0.021 0.329$\pm$0.006 0.186$\pm$0.005 0.593$\pm$0.022 0.741$\pm$0.063 0.350$\pm$0.009 0.222$\pm$0.008
0.498$\pm$0.006 0.482$\pm$0.006 0.234$\pm$0.005 0.107$\pm$0.001 0.437$\pm$0.006 0.397$\pm$0.005 0.286$\pm$0.005 0.167$\pm$0.004
$\downarrow$ 10.8% 21.1% 28.9% 42.5% 26.3% 46.4% 18.3% 24.8%
Appendix E Performance Comparison of CondTSF and Smoothing
We distill the dataset using the one-shot condensing ratio. The information on condensation is shown in Table.2. We conduct experiments on distilling dataset with CondTSF and a low pass filter respectively. We use DLinear[45] as the expert model. The result is shown in Table.11.
Results indicate that using CondTSF is significantly more effective than using a low pass filter to smooth the distilled data.
Table 11: Distill performance of different dataset condensation methods. Five synthetic datasets are distilled and the average and standard deviation are reported.
ExchangeRate Weather Electricity Traffic
MAE MSE MAE MSE MAE MSE MAE MSE
Random 0.783$\pm$0.090 1.070$\pm$0.246 0.530$\pm$0.084 0.647$\pm$0.159 0.840$\pm$0.017 1.102$\pm$0.031 0.854$\pm$0.018 1.350$\pm$0.043
Smooth 0.867$\pm$0.106 1.358$\pm$0.321 0.602$\pm$0.018 0.772$\pm$0.032 0.831$\pm$0.042 1.068$\pm$0.109 0.786$\pm$0.037 1.208$\pm$0.101
MTT+Smooth 0.620$\pm$0.024 0.636$\pm$0.062 0.501$\pm$0.012 0.527$\pm$0.028 0.788$\pm$0.027 0.963$\pm$0.061 0.836$\pm$0.022 1.291$\pm$0.057
MTT+CondTSF 0.195$\pm$0.007 0.061$\pm$0.004 0.326$\pm$0.009 0.284$\pm$0.007 0.391$\pm$0.003 0.284$\pm$0.004 0.494$\pm$0.022 0.579$\pm$0.037
ETTm1 ETTm2 ETTh1 ETTh2
MAE MSE MAE MSE MAE MSE MAE MSE
Random 0.728$\pm$0.033 0.993$\pm$0.082 0.695$\pm$0.011 0.889$\pm$0.030 0.756$\pm$0.035 1.059$\pm$0.083 0.749$\pm$0.037 1.013$\pm$0.089
Smooth 0.728$\pm$0.013 0.991$\pm$0.038 0.706$\pm$0.053 0.910$\pm$0.136 0.701$\pm$0.007 0.879$\pm$0.026 0.744$\pm$0.067 1.023$\pm$0.145
MTT+Smooth 0.656$\pm$0.014 0.919$\pm$0.040 0.619$\pm$0.029 0.684$\pm$0.062 0.680$\pm$0.012 0.874$\pm$0.044 0.704$\pm$0.053 0.910$\pm$0.136
MTT+CondTSF 0.491$\pm$0.004 0.502$\pm$0.008 0.347$\pm$0.028 0.202$\pm$0.028 0.532$\pm$0.014 0.580$\pm$0.029 0.329$\pm$0.003 0.205$\pm$0.002
Appendix F Ablation Study of CondTSF